OpenStudy (anonymous):

2 trains 40 miles apart head toward each other at 10mph each. A fly flying 15 miles an hour flies back and forth between them. How far foes the fly fly until its horrible demise?

OpenStudy (bahrom7893):

LOL funny question...

OpenStudy (bahrom7893):

hmm... okay so let's use logic: if we have two trains that both need to cover 20 miles at 10mph each.. since they are going to crash halfway..

OpenStudy (bahrom7893):

So it will take the first train: t = d/V = 20/10 = 2hrs to get to the middle.. which means that the other train will get there in two hours as well..

OpenStudy (bahrom7893):

At 15 miles per hour... the fly travels: d = V*t = 15 * 2 = 30 miles..

OpenStudy (bahrom7893):

The fly flies 30 miles until its horrible demise.